# On an approach to canonicalizing elliptic Feynman integrals

### v2

@@ Abstract
We develop a systematic approach to systematically constructing canonical bases for univariate elliptic Feynman integrals with multiple kinematic scales, which frequently arise in phenomenologically relevant scattering processes. The construction is performed in the Legendre normal form of elliptic curves, where the geometric properties of the curves are simple and explicit, and further kinematic singularities are present as marked points. The simplicity of the normal form facilitates the construction of universal canonical bases with an arbitrary number of marked points, which can be mapped to generic univariate elliptic integral families via an appropriate Möbius transformation. Therefore, our result can be readily applied to these cases straightforwardly. As a demonstration, we discuss the application of our method to several concrete examples, where we show that the $\varepsilon$-factorization of sub-sector dependence can also be simply done.
@@ 1 Introduction
The differential equations method Kotikov:1990kg ; Kotikov:1991pm ; Remiddi:1997ny ; Gehrmann:1999as , and specifically the canonical differential equations approach Henn:2013pwa , represents the state-of-the-art method for calculating MIs analytically. This approach leverages an $\varepsilon$-factorized form, allowing MIs to be expressed as Laurent expansions in the dimensional regulator $\varepsilon$, with Chen’s iterated integrals Chen:1977oja as coefficients. [^1] A canonical basis has other nice properties in addition to $\varepsilon$-factorization. The simplest iterated integrals found in Feynman integrals are multiple polylogarithms (MPLs) Goncharov:1998kja ; Goncharov:2001iea ; Vollinga:2004sn , which are iterated integrals over the genus-00 Riemann sphere with marked points, or say over moduli space ${\cal M}_{0,n}$ Bogner:2014mha . For MPL integral families, canonical differential equations can be achieved through rational and algebraic transformations alone, and several tools are available to assist in identifying such transformations Gituliar:2017vzm ; Prausa:2017ltv ; Meyer:2017joq ; Lee:2020zfb ; Dlapa:2020cwj . An alternative approach constructs canonical bases through ${\rm d}\log$-forms Henn:2020lye ; Chen:2020uyk ; Chen:2022lzr . [^2] However, not all integrals in ${\rm d}\log$-forms are polylogarithmic, see Duhr:2020gdd . MPLs exhibit properties that facilitate algorithmic extraction of critical information, aiding in the construction of canonical bases and enriching our understanding of their mathematical structure.
@@
However, as the complexity of the problem increases with additional loops and mass scales, the MPL functions space becomes insufficient to accommodate all integrals within an integral family (see, e.g., Bourjaily:2022bwx for a recent review). In these cases, elliptic integrals—the next class of iterated integrals—emerge, first noted in the context of the two-loop electron self-energy in QED Sabry:1962rge . Fully analytic results for this case were obtained only half a century later Honemann:2018mrb , and additional significant calculations involving more scales or higher loops are found in works such as Bogner:2019lfa ; Giroux:2024yxu ; Pogel:2022yat ; Pogel:2022ken ; Pogel:2022vat . Three-loop corrections have also been computed in recent studies Duhr:2024bzt . Elliptic Feynman integrals are iterated integrals over a genus-$1$ Riemann surface, or say torus, with marked points, equivalently, over moduli space ${\cal M}_{1,n}$ Weinzierl:2020kyq . Unlike MPLs, which are defined on the Riemann sphere with a unique shape, the genus-$1$ Riemann surface underlying elliptic Feynman integrals varies and can be parameterized by a modular variable, $\tau$. To express the $\varepsilon$-expansion of elliptic Feynman integrals, we need both the elliptic multiple polylogarithms (eMPLs) Levin:2007tto ; Brown:2011wfj ; Broedel:2017kkb and iterated integrals of modular forms or Eisenstein series Adams:2017ejb ; Broedel:2018iwv ; Duhr:2019rrs ; Weinzierl:2020kyq . It’s well-known that the underlying geometry for an integral can be made explicit with the analysis of leading singularities Cachazo:2008vp ; Arkani-Hamed:2010pyv . Specially, the underlying elliptic curve of an elliptic sector can be identified in (loop-by-loop) Baikov representations Baikov:1996rk ; Baikov:1996iu ; Frellesvig:2017aai ; Harley:2017qut through the maximal cut Frellesvig:2017aai ; Harley:2017qut ; Primo:2016ebd ; Bosma:2017ens ; Primo:2017ipr . Implementations for (loop-by-loop) Baikov representations are offered by Frellesvig:2017aai ; Jiang:2023qnl ; Jiang:2024eaj ; Frellesvig:2024ymq ; Correia:2025yao . Alternatively, we can study the elliptic curve by analyzing the periods of the elliptic curve through solutions to a second-order irreducible Picard-Fuchs operator Adams:2017tga ; Adams:2018bsn ; Adams:2018kez .
@@
It is well-known that the shapes of the elliptic curves depend only on the elliptic modulus or, equivalently, the modular lambda function $\lambda$. This is particularly evident in the so-called normal forms of elliptic curves, such as the Legendre normal form or the Jacobi normal form. In this paper, we employ this fact to systematize the construction of canonical bases for univariate elliptic Feynman integrals under the maximal cut. We use the term “canonical” here since it meets most of the important properties advocated by the literature Dlapa:2022wdu ; Frellesvig:2023iwr ; Duhr:2024uid . Our method utilizes the Legendre normal form of elliptic curves, where the geometric properties of the curves are simple and explicit. [^3] The Legendre normal form has also been employed in Broedel:2018rwm . Similarly, the Rosenhain normal form for the hyperelliptic case has been explored recently in Duhr:2024uid . We then construct the pre-canonical basis using Abelian differentials of three kinds and straightforwardly convert them to the canonical form with a universal transformation. This transformation is found by a method similar to Pogel:2022vat , and the result is also equivalent to Gorges:2023zgv . Due to the simplicity of the normal form, we are able to generically write down the canonical basis for a Legendre elliptic family with an arbitrary number of kinematic marked points. This generic canonical basis is the main new result of this paper. Finally, the Möbius transformation relating an arbitrary elliptic curve and its Legendre normal form can be written in a generic way as well. With such a transformation, our result can be directly applied to any univariate elliptic Feynman integral families.
To demonstrate our construction, we apply the method to several concrete examples: fully massive sunrise and non-planar double box diagrams, each with two kinds of mass configurations. We show that it is straightforward to derive an $\varepsilon$-factorized system of differential equations within the top sectors. To obtain a canonical basis for the entire integral family, we still need to address the non-$\varepsilon$-factorized dependence on the sub-sectors. Ideally, one would imagine that it is possible to construct the full integrand without any cut [^4] We have demonstrated an example with the next-to-maximal cut in Sec. 3.2., combining the top-sector construction in this work and the ${\rm d}\log$-form construction in Chen:2020uyk ; Chen:2022lzr . In this work, we simply convert the differential equations to the $\varepsilon$-factorized form through a linear transformation. We show in the examples that the nice properties of our basis help a lot in this process.
@@
This paper is organized as follows. In Sec. 2, We present the construction of the canonical basis under maximal cut, which involves the construction in the Legendre normal form and the transformation to an elliptic curve defined by an arbitrary degree-4 polynomial. We demonstrate the procedure with several examples in Sec. 3, where the transformation to a canonical basis for the full family without cut is discussed. The summary and outlook are provided in Sec. 4. There are a few appendices at the end of the paper. A different strategy vonManteuffel:2017hms ; Frellesvig:2021hkr for deriving an $\varepsilon$-factorized basis is presented in App. A. The construction using the Jacobi normal form of elliptic curves is discussed in App. B. Some lengthy expressions relevant to the examples in Sec. 3 are collected in App. C. Special relations between elliptic integrals are compiled in App. D.
@@ 2 Canonical basis under the maximal cut
In this section, we first briefly review some basics for elliptic Feynman integrals. Then, we present a general approach to construct a pre-canonical basis in elliptic sectors, which can be easily converted to a canonical basis under the maximal cut. We will first perform the construction in terms of the Legendre form $P_{L}(x)$ of the elliptic curve and then transform to the quartic family $P_{4}(z)$, which is commonly used in the Baikov representation with a suitable Möbius transformation. Finally, we construct the transformation to the canonical basis, which can be used directly in both the Legendre family and the quartic family.
2.1 Basics for elliptic Feynman integrals
Integral family
@@ 2.1 Basics for elliptic Feynman integrals
with $N=L(L+1)/2+LE$, $d=d_{0}-2\varepsilon$ ($d_{0}\in 2\mathbb{Z}^{+}$) and $\nu=\sum_{i=1}^{N}\nu_{i}$. A topology in the family is specified by the set of $D_{i}$ with the corresponding $\nu_{i}$ positive, and the rest of the propagators are irreducible scalar products (ISPs). Unless otherwise stated, we will consider integral families in $d=4-2\varepsilon$.
Elliptic curve
To identify the elliptic curve involved in an integral family, we resort to the maximal cut of the elliptic sector integrals in the (loop-by-loop) Baikov representation. Then the integral family under the cut can be expressed as
@@
where the multi-valued function $u$ (the twist in the language of twisted cohomology) is [^5] Here and in the following, we always consider the univariate elliptic family and suppress factors that only depend on kinematic variables in $u$.
$u=u(z)=[P_{4}(z)]^{-1/2}\prod_{i=1}^{n}(z-c_{i})^{-\beta_{i}\varepsilon}\,,$ (3)
where the branch points $c_{i}$’s are distinct and the corresponding $\beta_{i}$’s are integers. The single-valued differential $1$-form $\varphi=\varphi(z)$ may have poles at some of the branch points $c_{i}$’s. Note that there is an implicit branch point at infinity, which will be denoted as $c_{n+1}\equiv c_{\infty}\equiv\infty$. The elliptic curve is specified by the polynomial equation
$w^{2}=P_{4}(z)\equiv(z-c_{1})(z-c_{2})(z-c_{3})(z-c_{4})\,,$ (4)
which is only related to the four branch points $c_{1},c_{2},c_{3}$ and $c_{4}$.
The elliptic curve (4) can be represented in different ways, related by Möbius transformations $(z,w)\mapsto(x,y)$. The shape of the elliptic curve is characterized by the modular lambda $\lambda$,
@@
The Legendre normal form of the elliptic curve is given by
@@
with four branch points $10,\lambda,$ 1 and $\infty$. These four points are related to the previous four branch points, $c_{1}-c_{4}$, of the multi-valued function $w=\sqrt{P_{4}(z)}$ by a Möbius transformation. We will mainly use this form in the construction of the pre-canonical master integrals. In App. B, we will discuss the usage of a different form.
Elliptic integrals
We consider elliptic integrals of the form
$\int\frac{R(x)\,{\rm d}x}{\sqrt{P_{L}(x)}}\,,$ (7)
where $R(x)$ is a rational function. They can be converted to combinations of (3 kinds of) standard elliptic integrals by the variable change $x=\lambda t^{2}$:

### v3

@@ Abstract
We present generic expressions for the integrands of canonical bases under maximal cut in elliptic Feynman integral families with multiple kinematic scales. Such integrals frequently arise in phenomenologically relevant scattering processes. The derivation of our results starts from the Legendre normal form of elliptic curves, where the geometric properties of the curves are simple and explicit, and further kinematic singularities are presented as marked points. The simplicity of the normal form allows a straightforward construction of canonical bases with an arbitrary number of marked points. They can then be mapped into any univariate elliptic integral families via an appropriate Möbius transformation, leading to universal expressions for the integrands. As a demonstration, we discuss the application of our method to several concrete examples, including two new integral families whose canonical bases were not available in the literature. In several examples, we derive canonical bases for the full integral families without any cuts, demonstrating the simplicity of the sub-sector dependence of our canonical bases.
@@ 1 Introduction
The differential equations method Kotikov:1990kg ; Kotikov:1991pm ; Remiddi:1997ny ; Gehrmann:1999as , and specifically the canonical differential equations approach Henn:2013pwa , represents the state-of-the-art method for calculating MIs analytically. This approach leverages an $\varepsilon$-factorized form, allowing MIs to be expressed as Laurent expansions in the dimensional regulator $\varepsilon$, with Chen’s iterated integrals Chen:1977oja as coefficients. [^1] A canonical basis has other nice properties in addition to $\varepsilon$-factorization. The simplest iterated integrals found in Feynman integrals are multiple polylogarithms (MPLs) Goncharov:1998kja ; Goncharov:2001iea ; Vollinga:2004sn , which are iterated integrals over the genus-0 Riemann sphere with marked points, or say over moduli space ${\cal M}_{0,n}$ Bogner:2014mha . For MPL integral families, canonical differential equations can be achieved through rational and algebraic transformations alone, and several tools are available to assist in identifying such transformations Gituliar:2017vzm ; Prausa:2017ltv ; Meyer:2017joq ; Lee:2020zfb ; Dlapa:2020cwj . An alternative approach constructs canonical bases through ${\rm d}\log$-forms Henn:2020lye ; Chen:2020uyk ; Chen:2022lzr . [^2] However, not all integrals in ${\rm d}\log$-forms are polylogarithmic, see Duhr:2020gdd . Basis integrals derived this way automatically possess uniform transcendentality (UT) Henn:2013pwa , which is another key property of canonical bases in addition to $\varepsilon$-factorization. MPLs exhibit properties that facilitate algorithmic extraction of critical information, aiding in the construction of canonical bases and enriching our understanding of their mathematical structure.
@@
However, as the complexity of the problem increases with additional loops and mass scales, the MPL functions space becomes insufficient to accommodate all integrals within an integral family (see, e.g., Bourjaily:2022bwx for a recent review). In these cases, elliptic integrals—the next class of iterated integrals—emerge, first noted in the context of the two-loop electron self-energy in QED Sabry:1962rge . Fully analytic results for this case were obtained only half a century later Honemann:2018mrb , and additional significant calculations involving more scales or higher loops are found in works such as Bogner:2019lfa ; Giroux:2024yxu ; Pogel:2022yat ; Pogel:2022ken ; Pogel:2022vat . Three-loop corrections have also been computed in recent studies Duhr:2024bzt . Elliptic Feynman integrals are iterated integrals over a genus-$1$ Riemann surface, or say torus, with marked points, equivalently, over moduli space ${\cal M}_{1,n}$ Weinzierl:2020kyq . Unlike MPLs, which are defined on the Riemann sphere with a unique shape, the genus-$1$ Riemann surface underlying elliptic Feynman integrals varies and can be parameterized by a modular variable, $\tau$. To express the $\varepsilon$-expansion of elliptic Feynman integrals, we need both the elliptic multiple polylogarithms (eMPLs) Levin:2007tto ; Brown:2011wfj ; Broedel:2017kkb and iterated integrals of modular forms or Eisenstein series Adams:2017ejb ; Broedel:2018iwv ; Duhr:2019rrs ; Weinzierl:2020kyq . It’s well-known that the underlying geometry for an integral can be made explicit with the analysis of leading singularities Cachazo:2008vp ; Arkani-Hamed:2010pyv . Specifically, the underlying elliptic curve of an elliptic sector can be identified in (loop-by-loop) Baikov representations Baikov:1996rk ; Baikov:1996iu ; Frellesvig:2017aai ; Harley:2017qut through the maximal cut Frellesvig:2017aai ; Harley:2017qut ; Primo:2016ebd ; Bosma:2017ens ; Primo:2017ipr . Implementations for (loop-by-loop) Baikov representations are offered by Frellesvig:2017aai ; Jiang:2023qnl ; Jiang:2024eaj ; Frellesvig:2024ymq ; Correia:2025yao . Alternatively, we can study the elliptic curve by analyzing the periods of the elliptic curve through solutions to a second-order irreducible Picard-Fuchs operator Adams:2017tga ; Adams:2018bsn ; Adams:2018kez .
@@
In this paper, we present universal expressions for the integrands of canonical bases in any univariate elliptic integral families under the maximal cut. The derivation of the integrands utilizes the Legendre normal form [^3] The Legendre normal form has also been employed in Broedel:2018rwm . Similarly, the Rosenhain normal form for the hyperelliptic case has been explored recently in Duhr:2024uid . of elliptic curves, where the geometric information is evident through the elliptic modular lambda function $\lambda$. The results are then mapped to a generic elliptic family via a Möbius transformation. We use the term “canonical” here since they exhibit many of the important properties of canonical bases advocated by the literature Dlapa:2022wdu ; Frellesvig:2023iwr ; Duhr:2024uid , besides $\varepsilon$-factorization. In particular, they approach $\mathrm{d}\log$ forms at the degenerate limit of the elliptic curves, where the differential equations also exhibit a simple pole behavior.
Our results are at the integrand level, and do not require further information from the differential equations within the top sector. We apply our method to several concrete examples, including the unequal-mass sunrise family and non-planar double-box families. We further consider two phenomenologically relevant integral families, and provide their canonical bases for the first time. In several examples, we obtain canonical bases for the full integral families without any cuts. This benefits from the simplicity of the sub-sector dependence of our pre-canonical bases. To further simplify the sub-sector dependence, we make preliminary attempts for multivariate constructions under the next-to-maximal cut, where we show that a large portion of the differential equations can be automatically rendered $\varepsilon$-factorized.
@@
This paper is organized as follows. In Sec. 2, we present the main results, the universal expressions for the canonical bases, and their derivation. We apply the results to several examples in Sec. 3. The summary and outlook are provided in Sec. 4. Several appendices are included at the end of the paper. A different strategy vonManteuffel:2017hms ; Frellesvig:2021hkr for deriving an $\varepsilon$-factorized basis is presented in App. A. The construction using the Jacobi normal form of elliptic curves is discussed in App. B. Some lengthy expressions relevant to the example in Sec. 3.4 are collected in App. C. Special relations between elliptic integrals are compiled in App. D. The canonical bases for all examples are offered in the attached ancillary files.
@@ 2 Canonical basis under the maximal cut
2.1 Integral family and canonical basis
@@
with $N=L(L+1)/2+LE$, $d=d_{0}-2\varepsilon$ ($d_{0}\in 2\mathbb{Z}^{+}$) and $\nu=\sum_{i=1}^{N}\nu_{i}$. A sector in the family is specified by the set of $D_{i}$ with the corresponding $\nu_{i}$ positive, and the rest of the propagators are irreducible scalar products (ISPs). Unless otherwise stated, we will consider integral families in $d=4-2\varepsilon$.
We consider an elliptic sector within this integral family in the (loop-by-loop) Baikov representation Frellesvig:2017aai ; Jiang:2023qnl ; Jiang:2024eaj ; Frellesvig:2024ymq ; Correia:2025yao . We assume that there is only one ISP variable $z$ in this representation, such that after applying maximal cut, the integrals in this sector can be expressed as
@@
where the multi-valued function $u$ (the twist in the language of twisted cohomology) is [^4] Here and in the following, we always suppress factors that only depend on kinematic variables in $u$.
$u=u(z)=[P_{4}(z)]^{-1/2}\prod_{i=1}^{n}(z-c_{i})^{-\beta_{i}\varepsilon}\,,\quad P_{4}(z)\equiv(z-c_{1})(z-c_{2})(z-c_{3})(z-c_{4})\,$, (3)
where the branch points $c_{i}$’s are distinct and the corresponding $\beta_{i}$’s are integers. The single-valued differential $1$-form $\varphi$ may have poles at some of the branch points $c_{i}$’s. Note that there is an implicit branch point at infinity, which will be denoted as $c_{n+1}\equiv c_{\infty}\equiv\infty$. For convenience, we assume $c_{1}<c_{2}<c_{3}<c_{4}$, and introduce the reduced twist
$\bar{u}(z)=\prod_{i=1}^{n}(z-c_{i})^{-\beta_{i}\varepsilon}\,$, (4)
so that the integrals can be represented as
$I=\int_{\cal C}\bar{u}(z)\frac{\varphi}{\sqrt{P_{4}(z)}}\equiv\int_{\cal C}\bar{u}(z)\,\phi\,.$ (5)
The degree-4 polynomial $P_{4}(z)$ defines an elliptic curve $w^{2}=P_{4}(z)$. The shape of the elliptic curve is characterized by the modular lambda $\lambda$,
@@
The elliptic curve can be transformed into the Legendre normal form
@@
by a Möbius transformation
$z\mapsto x=T(z)=\frac{(z-c_{2})c_{14}}{(z-c_{1})c_{24}}\,.$ (8)
Under this Möbius transformation, the branch points are mapped as
$c_{1}$ $\mapsto e_{1}=\infty\,,\quad c_{2}\mapsto e_{2}=0\,,\quad c_{3}\mapsto e_{3}=\lambda\,,\quad c_{4}\mapsto e_{4}=1\,$,
$c_{i}$ $\mapsto e_{i}=\frac{c_{14}c_{2i}}{c_{1i}c_{24}}\,,\;(i=5,\cdots,n)\,,\quad c_{n+1}=\infty\mapsto e_{n+1}\equiv e_{\infty}=\frac{c_{14}}{c_{24}}\,.$ (9)
We can now state the explicit form for the integrands $\phi_{i}^{\text{can}}$ of the canonical basis $I^{\text{can}}_{i}$ for $i=1,\cdots,n-1$:
$\phi_{1}^{\rm can}$ $=\frac{1}{\varpi_{1}(\lambda)}\frac{\sqrt{c_{13}c_{24}}}{\sqrt{P_{4}(z)}}{\rm d}z\,$, (10a)
$\phi_{i-3}^{\rm can}$ $=\left(\frac{1}{z-c_{i}}-\frac{1}{c_{1i}}\right)\frac{\sqrt{P_{4}(c_{i})}}{\sqrt{P_{4}(z)}}{\rm d}z-\frac{\vartheta_{1}(e_{i},\lambda)}{\varpi_{1}(\lambda)}\frac{\sqrt{c_{13}c_{24}}}{\sqrt{P_{4}(z)}}{\rm d}z\,,\quad(i=5,\cdots,n)\,$, (10b)
$\phi_{n-2}^{\rm can}$ $=-\frac{z-c_{1}}{\sqrt{P_{4}(z)}}{\rm d}z-\frac{\vartheta_{1}(e_{n+1},\lambda)}{\varpi_{1}(\lambda)}\frac{\sqrt{c_{13}c_{24}}}{\sqrt{P_{4}(z)}}{\rm d}z\,$, (10c)
$\phi_{n-1}^{\rm can}$ $=-\frac{1+2\beta_{1}\varepsilon}{8\varepsilon}\varpi_{1}(\lambda)\frac{c_{41}(z-c_{2})}{c_{42}(z-c_{1})}\frac{\sqrt{c_{13}c_{24}}}{\sqrt{P_{4}(z)}}{\rm d}z-\frac{1}{4}\sum_{i=5}^{n}\beta_{i}\vartheta_{1}(e_{i},\lambda)\left(\frac{1}{z-c_{i}}-\frac{1}{c_{1i}}\right)\frac{\sqrt{P_{4}(c_{i})}}{\sqrt{P_{4}(z)}}{\rm d}z$
$+\frac{1}{4}\beta_{n+1}\vartheta_{1}(e_{n+1},\lambda)\frac{z-c_{1}}{\sqrt{P_{4}(z)}}{\rm d}z-\frac{1}{8\varepsilon\varpi_{1}(\lambda)}\Bigg\{\lambda\varpi_{1}(\lambda)[2(1-\lambda)\varpi_{1}^{\prime}(\lambda)-\varpi_{1}(\lambda)]$
$\hskip 100.00015pt+\varepsilon\sum_{i=2}^{n+1}\beta_{i}(1+\lambda-e_{i})\varpi_{1}(\lambda)^{2}-\varepsilon\sum_{i=5}^{n+1}\beta_{i}\vartheta_{1}(e_{i},\lambda)^{2}\Bigg\}\frac{\sqrt{c_{13}c_{24}}}{\sqrt{P_{4}(z)}}{\rm d}z\,$, (10d)
where $\varpi_{1}$ and $\vartheta_{1}$ are related to the complete elliptic integrals of the first and the third kinds, respectively:
$\varpi_{1}(\lambda)$ $\equiv\frac{2i}{\pi}\int\limits_{\lambda}^{1}\frac{{\rm d}x}{\sqrt{P_{L}(x)}}=\frac{4}{\pi}K(1-\lambda)\,$, (11a)
$\vartheta_{1}(e_{i},\lambda)$ $\equiv-\frac{2i\sqrt{P_{L}(e_{i})}}{\pi e_{i}}\int\limits_{\lambda}^{1}\frac{1}{1-x/e_{i}}\,\frac{{\rm d}x}{\sqrt{P_{L}(x)}}=-\frac{4\sqrt{P_{L}(e_{i})}}{\pi e_{i}}\left[K(1-\lambda)+\frac{\lambda}{e_{i}-\lambda}\Pi\left(\frac{e_{i}(1-\lambda)}{e_{i}-\lambda},1-\lambda\right)\right]\,.$ (11b)
In later derivations of Eq. (10), we will show that the differential equations of $\{I_{i}^{\text{can}}\}$ are $\varepsilon$-factorized under maximal cut. However, $\varepsilon$-factorization is only a necessary condition for a canonical basis. In the polylogarithmic case, we would impose further requirements besides $\varepsilon$-factorization, to ensure that the canonical integrals evaluate to pure functions. One such requirement is that the integrals should degenerate to UT constants at certain kinematic boundary points. In the elliptic case, one may expect similar properties from a canonical basis Frellesvig:2023iwr . In particular, in the kinematic limits where the elliptic curves degenerate to Riemann spheres, one would expect that the canonical integrals degenerate to UT polylogarithmic functions.
With this in mind, it is interesting to investigate the behavior of the basis (10) in the limit $c_{1}\to c_{2}$, where $\lambda\to 1$ and the elliptic curve degenerates. One can show that the integrands $\phi_{i}^{\text{can}}$ degenerate to ${\rm d}\log$-forms in the limit $c_{1}\to c_{2}$:
$\lim_{c_{1}\to c_{2}}\phi_{1}^{\rm can}$ $=-\frac{1}{2}\,{\rm d}\log\left(\frac{1+\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}{1-\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}\right)\,$, (12a)
$\lim_{c_{1}\to c_{2}}\phi_{i-3}^{\rm can}$ $={\rm d}\log\left(\frac{1+\sqrt{\frac{c_{i4}(z-c_{3})}{c_{i3}(z-c_{4})}}}{1-\sqrt{\frac{c_{i4}(z-c_{3})}{c_{i3}(z-c_{4})}}}\right)-{\rm d}\log\left(\frac{1+\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}{1-\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}\right)\,$, (12b)
$\lim_{c_{1}\to c_{2}}\phi_{n-2}^{\rm can}$ $={\rm d}\log\left(\frac{1+\sqrt{\frac{z-c_{3}}{z-c_{4}}}}{1-\sqrt{\frac{z-c_{3}}{z-c_{4}}}}\right)-{\rm d}\log\left(\frac{1+\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}{1-\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}\right)\,$, (12c)
$\lim_{c_{1}\to c_{2}}\phi_{n-1}^{\rm can}$ $=\frac{1}{4}(2\beta_{1}+2\beta_{2}+\beta_{3}+\beta_{4})\,{\rm d}\log\left(\frac{1+\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}{1-\sqrt{\frac{c_{41}(z-c_{3})}{c_{31}(z-c_{4})}}}\right)+\frac{1}{2}\sum_{i=5}^{n+1}\beta_{i}\,{\rm d}\log\left(\frac{1+\sqrt{\frac{c_{i4}(z-c_{3})}{c_{i3}(z-c_{4})}}}{1-\sqrt{\frac{c_{i4}(z-c_{3})}{c_{i3}(z-c_{4})}}}\right)\,.$ (12d)
Since ${\rm d}\log$-form integrands lead to UT integrals, we see that the basis in Eq. (10) is indeed a natural generalization of the polylogarithmic canonical basis to elliptic cases.
The explicit expressions of the canonical basis (10), together with the Möbius transformation (8), are the main results of this work. They are generic enough to be directly applied to any elliptic integral family with an arbitrary number of branch points. In the following subsections, we show how the above results are derived. In the next section, we apply the above construction to several examples, presenting novel results not available in the literature, to demonstrate the power of our method.
2.2 The construction of the pre-canonical basis with the Legendre normal form
To derive the canonical basis (10), we start from a pre-canonical basis for the integrand after the Möbius transformation (8). The transformed $u$ function is given by
$u_{L}(x)=[P_{L}(x)]^{-1/2}\prod_{i=2}^{n+1}(x-e_{i})^{-\beta_{i}\varepsilon}\equiv[P_{L}(x)]^{-1/2}\,\bar{u}_{L}(x)\,$, (13)
where
$\beta_{n+1}=-\sum_{i=1}^{n}\beta_{i}\,.$ (14)
The integrals in this sector under maximal cut then have the general form
$I=\int_{\cal C}u_{L}(x)\,\varphi=\int_{\cal C}\bar{u}_{L}(x)\,\phi\,$, (15)
where as before, we have absorbed the factor $P_{L}(x)]^{-1/2}$ into $\phi$ for later convenience.
To find suitable forms of $\phi$, we take hints from the three kinds of standard elliptic integrals. The incomplete elliptic integral of the first kind is defined as
where the differential is holomorphic. For the second equal-sign in the above formula, we have used